# Hilbert space methods for approximating multi-output latent variable Gaussian processes

### v1

@@ Abstract
Gaussian processes are a powerful class of non-linear models, but have limited applicability for larger datasets due to their high computational complexity. In such cases, approximate methods are required, for example, the recently developed class of Hilbert space Gaussian processes. They have been shown to drastically reduce computation time while retaining most of the favourable properties of exact Gaussian processes. However, Hilbert space approximations have so far only been developed for uni-dimensional outputs and manifest (known) inputs. To this end, we generalise Hilbert space methods to multi-output and latent input settings. Through extensive simulations, we show that the developed approximate Gaussian processes are indeed not only faster, but also provides similar or even better uncertainty calibration and accuracy of latent variable estimates compared to exact Gaussian processes. While not necessarily faster than alternative Gaussian process approximations, our new models provide better calibration and estimation accuracy, thus striking an excellent balance between trustworthiness and speed. We additionally validate our findings in a real world case study from single cell biology.
@@ 1 Introduction
Gaussian processes (GPs) are a popular class of non-linear models well-known for their flexibility in modelling a wide range of data scenarios (Williams and Rasmussen, , 1995). They are also well-known for their steep computational complexity, scaling cubically with the sample size. Thus, exact GPs cannot be used to model large datasets within feasible time. These scalability issues are even more pronounced for extensions of standard GP models. Here, we consider a combination of GPs for latent variable estimation (Lawrence, , 2003, 2005) and GPs with multi-dimensional outputs (Cressie, , 1993; Teh et al., , 2005). Multi-output GPs are tasked to jointly model several, potentially correlated outcome variables, each with their own GP. In turn, estimating latent variables via GPs not only increases the number of estimable parameters drastically, but also requires multiple (potentially many) outputs to achieve sufficient estimation accuracy. Multi-output GPs for latent variable estimation have important applications, for example, in single-cell biology where estimating latent cell ordering from multi-dimensional RNA gene sequencing data is crucial to understand the dynamics of the biological process (Haque et al., , 2017). Yet, when using exact GPs for this purpose, only small datasets can be analysed, which strongly limits their practical applicability.
@@
An alternative approach is to approximate the covariance function through its spectral decomposition computed from a finite set of representative basis functions. The latter method falls under the category of Hilbert space approximate GPs (HSGPs; Solin and Särkkä, , 2020; Riutort-Mayol et al., , 2022). By exploiting the spectral representation of a stationary covariance function, the computational complexity of HSGPs scales linearly with both sample size and the number of basis functions. What is more, they come with powerful diagnostics that indicate whether the chosen number of basis functions was sufficient for an accurate GP approximation given the data at hand (Riutort-Mayol et al., , 2022). HSGPs are typically estimated via Markov-chain Monte Carlo (MCMC), which is slower than VI but produces posterior approximations of much higher quality. The increased, diagnosable accuracy of HSGPs combined with their efficiency has a lot of potential for latent variable estimation. However, HSGPs have so far only been developed for single outputs and manifest (known) inputs.
In this paper, we generalise HSGPs to multiple outputs and latent variable inputs. We validate our new HSGPs against the corresponding exact GPs (where feasible) as well as against variational GPs based on inducing points. We design elaborate simulation studies to investigate the properties of HSGPs in terms of latent variable estimation accuracy, uncertainty calibration and model convergence stability. Through our simulation studies, we provide a thorough comparison of HSGPs with exact GPs and variational GPs before applying them on a real-world single-cell RNA sequencing data.
@@ 2 Related work
In terms of real-world applications, we highlight the field of single-cell biology. There, multi-output structures naturally arise in a plethora of research problems (Haque et al., , 2017) and latent GPs are common to model cellular ordering (Hensman et al., 2013b, ; Campbell and Yau, , 2015). Later works involve estimating the cellular ordering (Reid and Wernisch, , 2016; Campbell and Yau, , 2018) along with inference on different cellular pathway dynamics. A direct combination of the above multi-output latent GPs for single-cell RNA sequencing data is presented in Ahmed et al., (2019), which used the combination of inducing points and VI methods for estimating cellular ordering. As a real-world case study, we showcase our models for a single-cell data depicting cyclical cellular pathway dynamics (Mahdessian et al., , 2021).
@@ 3 Methods
We initially discuss the structure of exact GPs and HSGPs for univariate output $y$ and covariate (known) with paired samples $(y_{i},x_{i})$ for $i={1,2,...,N}$ where $N$ is the sample size. Then, we generalise the Hilbert space framework for multi-dimensional outputs and latent inputs before discussing methods of inference and alternate GP approximation methods.
@@ 3.1 Gaussian processes
$y_{i}\sim\mathcal{N}(f(x_{i}),\sigma^{2}).$ (2)
For $i\neq j$, $\text{Cov}(y_{i},y_{j})=k(x_{i},x_{j})$ such that $i,j={1,2,...,N}$. When $i=j$, $\text{Cov}(y_{i},y_{j})=\text{Var}(y_{i})=k(x_{i},x_{j})+\sigma^{2}$. In terms of $k$, the Matern class of stationary covariance functions is a highly common choice (Williams and Rasmussen, , 1995; Rasmussen and Williams, , 2006). Among them, we specifically consider Squared Exponential (SE), Matern 3/2 and 5/2, which are defined as:
@@
where $r=|x-x^{\prime}|$ is the distance between two arbitrary inputs, $\alpha$ is the marginal standard deviation (SD) and $\rho$ is the length scale parameter. The above covariance functions vary in terms of their implied functional smoothness, with the SE covariance function being the most and the Matern 3/2 being the least smooth. We use a constant mean function $\mu$ for all of our exact GPs and HSGPs discussed in this paper.
@@ 3.2 Hilbert space approximations
where $\omega\in\mathbf{R}$ is the input in the frequency domain, and $\rho$ and $\alpha^{2}$ are the covariance function hyperparameters. The spectral densities can be derived for the general Matern class of covariance functions (Riutort-Mayol et al., , 2022), but for this paper, we specifically focus on the above special cases.
@@
$-\nabla^{2}\phi_{j}(x)$ $=\lambda_{j}\phi_{j}(x),\quad x\in\Omega$ (6)
$\phi_{j}(x)$ $=0,\quad x\notin\Omega$
@@ 3.3.1 Multi-output HSGPs
$y_{di}\sim\mathcal{N}(f^{*}_{d}(x_{i}),\sigma_{d}^{2})$ (15)
@@ 3.3.2 Latent variable HSGPs
Within latent-variable GPs, the inputs $x$ are considered as unobserved and are treated similar to other estimable parameters. To that end, we specify an observed quantity $\tilde{x}$ that guides the estimation of the latent inputs $x$. From a Bayesian perspective, $\tilde{x}$ acts as a prior for the latent $x$ to be then further refined by the GPs learning from $y$. Specifically, we consider the scenario where $\tilde{x}$ is a noisy measurement of $x$. If we assume that the measurements $\tilde{x}$ are Gaussian with known measurement SD $s$, we can write for each observation $i$:
@@ 3.4 Bayesian inference
$p(y,x,\theta)=p(x)\prod_{d}^{D}p(y_{d}\mid x,\theta_{d},)\,\,p(\theta_{d}).$ (18)
where $p(y_{d}\mid x,\theta_{d})$ denotes the GP-based likelihood for a single output dimension and $p(x)$ denotes the prior for the latent $x$ implied by the measurement model in Eq. (16). The details of prior specifications used in our experiments are further discussed in Section 4. Using Bayes’ rule, we obtain the joint posterior over $x$ and $\theta$ as
@@ 4 Simulation Study
For real-world data, we lack the ground truth values of latent variables. In contrast, in a simulated setting, we have full control over the data generating process including the true latent values. Thus, it is crucial to investigate the statistical properties of latent variable models using simulated data. To this end, we investigate exact GPs, HSGPs and VIGPs under various simulation scenarios exploring their behaviour in terms of model convergence, model calibration, and estimation accuracy. The data generating conditions are inspired by Mukherjee et al., (2025).
@@ 4.1 Data generating scenarios
In our simulations, we consider two version of this process, one with low oscillations (high true $\rho_{d}$ values) and one with higher oscillations (low true $\rho_{d}$ values). These two periodic data scenarios mimic the challenges of modelling starkly different functional smoothness with the higher oscillations being the more challenging one. The hyperparameters of the periodic process closely resemble that of the GP data scenarios and are thus named the same for simplicity. The sampling distributions are also similarly specified as mentioned previously. For the periodic data with higher oscillation, we specify $\rho\sim\text{Normal}^{+}(0.5,0.05^{2})$ as compared to the lower oscillation case where $\rho\sim\text{Normal}^{+}(1,0.25^{2})$.
@@ 4.2 Model specifications
We fit exact GPs, HSGPs, and VIGPs for all simulation scenarios, with the exception of omitting exact GPs for sample sizes higher than $N=20$ due to the scalability issues discussed earlier (see Section 1). For the HSGPs, we select the boundary conditions $L$ by multiplying a scaler adjustment $c=1.25$ to the range of the input prior $\tilde{x}$. This way, we ensure that we have a compact cover $[-L,L]$ that prevents input values $x$ to be near the boundaries. By doing so, we avoid model convergence issues completely (see Section 4.3). We select the minimum number of basis functions $M_{\rm min}$ loosely based on the suggestions of Riutort-Mayol et al., (2022). The choices for $M_{\rm min}$ depends on the choice of covariance function to accurately approximate the degree of non-linearity. Thus, For the SE covariance function,
@@
Both exact GPs and HSGPs are implemented in Stan (Stan Development Team, , 2024) and fitted with a single MCMC chain of 2000 iterations of which the first 1000 are discarded as warm-up. As shown in Mukherjee et al., (2025), model convergence is similar for multiple chains when fitting latent variable GPs. Thus, we run only a single chain per model to reduce overall computation times. The simulation studies are conducted on 50 vCPUs (Intel(R) Xeon(R) Gold 6230R CPU @ 2.10 GHz) with 720 GB work memory. The average runtime per dataset for the HSGPs with $N=200$ and $D=20$, the most computationally expensive case, was approximately 28 minutes for the SE model, 42 minutes for the Matern 3/2 model, and 34 minutes for the Matern 5/2 model. In the same scenarios, the VIGPs took approximately 2 minutes regardless of the covariance function. The exact GPs (just for $N=20$ and $D=20$) took 48 mins for the SE model, 1.6 hours for the Matern 3/2 and 2.4 hours for the Matern 5/2 model. The HSGPs were significantly faster than the exact GPs but slower than VIGPs. The slower speed of HSGPs compared to VIGPs is counter-balanced by highly consistent model calibrations, superior estimation accuracy for latent variables, and overall much more stable model performance, as shown below.
@@ 4.3 Model convergence
In Figure 1, we show $\widehat{R}$, Bulk-ESS, and Tail-ESS for the latent $x$ and GP hyperparameters of both exact GPs and HSGPs with SE covariance function. While the exact GPs reach and exceed the relaxed $\widehat{R}$ threshold of 1.1 for some simulated datasets, the HSGPs consistently satisfy the much stricter 1.01 threshold of model convergence. HSGPs subsequently also have much higher Bulk and Tail-ESS. Overall, based on the diagnostics, HSGPs show much more consistent and stable convergence as compared to exact GPs. We provide convergence diagnostic figures for the other simulation scenarios in the Appendix C (see Figures 11-15) where we see similar results.
@@ 4.4 Testing model calibration
To that end, briefly, SBC aims to test the goodness of posterior approximations by exploiting self-consistency properties of Bayesian models. This results in a rank statistic that is uniformly distributed under the assumption of a well calibrated model (Modrák et al., , 2023), which can subsequently be checked both graphically and numerically. As a graphical test, we plot the empirical cumulative distribution function (ECDF) of the ranks along with their 95% confidence regions under the assumption of uniformity (Säilynoja et al., , 2022). ECDFs lying outside of their confidence region indicate miscalibrations. These graphical tests also yield a test statistic (the log $\gamma$ score; see Säilynoja et al., , 2022), which we further analyse numerically. More details on SBC is provided in Appendix A.
Figure 3: Squared exponential scenario: log $\gamma$ scores offset by the 95% confidence threshold for all the fitted models. The behaviour of scores across true latent $x$ values are shown in the right-hand panel. The blue dashed line denotes the threshold to reject uniformity. The HSGP($M$) shows the HSGPs with their corresponding number of basis functions.
@@
To streamline calibration checking across simulation scenarios, we analyse the log $\gamma$ scores using a multi-level model (Bürkner, , 2017) described in Appendix A. In Figure 3, we show the predicted log $\gamma$ scores of the different GP models under the SE simulation scenario.
Figure 4: Squared exponential scenario (highly varying $\rho$): log $\gamma$ scores offset by the 95% confidence threshold for all the fitted models. The behaviour of scores across true latent $x$ values are shown in the right-hand panel. The blue dashed line denotes the threshold to reject uniformity. The HSGP($M$) shows the HSGPs with their corresponding number of basis functions.
For all the choices of $N$ and $D$, the latent variable estimates from HSGPs are consistently well calibrated. By comparison, the exact GPs fail the calibration test for $D=20$. The VIGPs shows miscalibrations especially for higher output dimension cases throughout all the choices of $N$. Additionally, exact GPs exhibit worse calibrations (decreasing log $\gamma$ scores) when estimating larger values of the latent inputs. This issue directly causes problems in estimating latent values closer to the boundary. VIGPs show similar behaviour for lower sample sizes $N=20$ although not as strong as exact GPs. The HSGPs are able to overcome and rectify this behaviour. For the special scenario of SE data with highly variable length-scales as shown in Figure 4, VIGPs and HSGPs with small number of basis function struggle to pass the calibration tests, especially for $N=200$. The miscalibrations are however alleviated with a higher number of basis functions in HSGPs. The log $\gamma$ results for the Matern 3/2 and 5/2 simulation scenarios (Figures 16 and 17) are presented in Appendix D.

### v2

@@ Abstract
Gaussian processes are a powerful class of non-linear models, but have limited applicability for larger datasets due to their high computational complexity. In such cases, approximate methods are required, for example, the recently developed class of Hilbert space Gaussian processes. They have been shown to drastically reduce computation time while retaining most of the favorable properties of exact Gaussian processes. However, Hilbert space approximations have so far only been developed for uni-dimensional outputs and manifest (known) inputs. To this end, we generalize Hilbert space methods to multi-output and latent input settings. Through extensive simulations, we show that the developed approximate Gaussian processes are indeed not only faster, but also provide similar or even better uncertainty calibration and accuracy of latent variable estimates compared to exact Gaussian processes. While not necessarily faster than alternative Gaussian process approximations, our new models provide better calibration and estimation accuracy, thus striking an excellent balance between trustworthiness and speed. We additionally validate our findings in a real world case study from single cell biology.
@@ 1 Introduction
Gaussian processes (GPs) are a popular class of non-linear models well-known for their flexibility in modeling a wide range of data scenarios (Williams and Rasmussen, , 1995). They are also well-known for their steep computational complexity, scaling cubically with the sample size. Thus, exact GPs cannot be used to model large datasets within feasible time. These scalability issues are even more pronounced for extensions of standard GP models. Here, we consider a combination of GPs for latent variable estimation (Lawrence, , 2003, 2005) and GPs with multi-dimensional outputs (Cressie, , 1993; Teh et al., , 2005). Multi-output GPs are tasked to jointly model several, potentially correlated outcome variables, each with their own GP. In turn, estimating latent variables via GPs not only increases the number of estimable parameters drastically, but also requires multiple (potentially many) outputs to achieve sufficient estimation accuracy. Multi-output GPs for latent variable estimation have important applications, for example, in single-cell biology where estimating latent cell orderings from multi-dimensional RNA gene sequencing data is crucial to understand the dynamics of the biological process (Haque et al., , 2017). Yet, when using exact GPs for this purpose, only small datasets can be analyzed, which strongly limits their practical applicability.
@@
An alternative approach is to approximate the covariance function through its spectral decomposition computed from a finite set of representative basis functions. The latter method falls under the category of Hilbert space approximate GPs (HSGPs; Solin and Särkkä, , 2020; Riutort-Mayol et al., , 2022). By exploiting the spectral representation of a stationary covariance function, the computational complexity of HSGPs scales linearly with both sample size and the number of basis functions. What is more, they come with powerful diagnostics that indicate whether the chosen number of basis functions was sufficient for an accurate GP approximation given the data at hand (Riutort-Mayol et al., , 2022). HSGPs are typically estimated via Markov-chain Monte Carlo (MCMC), which is slower than VI but produces posterior approximations of much higher quality. The increased, diagnosable accuracy of HSGPs combined with their efficiency has a lot of potential for latent variable estimation. However, HSGPs have so far only been developed for single outputs and manifest (known) inputs. We develop extensions for the HSGP framework to address both latent variable inputs and multi-dimensional outputs and compare their benefits over exact GPs as well as other GP approximation methods.
1.1 Overview of Contributions
  • We generalize HSGPs for latent variable inputs and multi-dimensional outputs.
  • We validate our new HSGPs against the corresponding exact GPs (where feasible) as well as against variational GPs based on inducing points approximation.
  • We perform extensive simulation studies to investigate the statistical properties of HSGPs in terms of latent variable estimation accuracy, uncertainty calibration, and model convergence.
  • We demonstrate the applicability of our HSGPs on real-world single-cell RNA sequencing data, whose analysis is infeasible with similarly specified exact GPs.
@@ 2 Related work
In terms of real-world applications, we highlight the field of single-cell biology. There, multi-output structures naturally arise in a plethora of research problems (Haque et al., , 2017) and latent GPs are common to model cellular ordering (Hensman et al., 2013b, ; Campbell and Yau, , 2015). Later works involve estimating the cellular ordering (Reid and Wernisch, , 2016; Campbell and Yau, , 2018) along with inference on different cellular pathway dynamics. A direct combination of the above multi-output latent GPs for single-cell RNA sequencing data is presented in Ahmed et al., (2019), which used the combination of inducing points and VI methods for estimating cellular ordering. As a real-world case study, we showcase our models for a single-cell data of the cell cycle (Mahdessian et al., , 2021).
@@ 3 Methods
We initially discuss the structure of exact GPs and HSGPs for univariate output $y$ and covariate (known) with paired samples $(y_{i},x_{i})$ for i=1,2,,Ni={1,2,...,N} where $N$ is the sample size. Then, we generalize the Hilbert space framework for multi-dimensional outputs and latent inputs before discussing methods of inference and alternate GP approximation methods.
@@ 3.1 Gaussian processes
$y_{i}\mid f\sim\mathcal{N}(f(x_{i}),\sigma^{2}).$ (2)
For $i\neq j$, $\text{Cov}(y_{i},y_{j})=k(x_{i},x_{j})$ such that i,j=1,2,,Ni,j={1,2,...,N}. When $i=j$, $\text{Cov}(y_{i},y_{j})=\text{Var}(y_{i})=k(x_{i},x_{j})+\sigma^{2}$. In terms of $k$, the Matern class of stationary covariance functions is a highly common choice (Williams and Rasmussen, , 1995; Rasmussen and Williams, , 2006). Among them, we specifically consider Squared Exponential (SE), Matern 3/2 and 5/2, which are defined as:
@@
where $r=|x-x^{\prime}|$ is the distance between two arbitrary inputs, $\alpha>0$ is the marginal standard deviation (SD) and $\rho>0$ is the length scale parameter. The above covariance functions vary in terms of their implied functional smoothness, with the SE covariance function being the most and the Matern 3/2 being the least smooth. We use a constant mean function $\mu$ for all of our exact GPs and HSGPs discussed in this paper.
@@ 3.2 Hilbert space approximations
where $\omega\in\mathbb{R}^{N}$ is the input in the frequency domain, and $\rho>0$ and $\alpha>0$ are the covariance function hyperparameters. The spectral densities can be derived for the general Matern class of covariance functions (Riutort-Mayol et al., , 2022), but for this paper, we specifically focus on the above special cases.
@@
$-\nabla^{2}\phi_{j}(x)$ $=\lambda_{j}\phi_{j}(x),\qquad x\in\Omega$ (6)
$\phi_{j}(x)$ $=0,\qquad x\notin\Omega$
@@ 3.3.1 Multi-output HSGPs
$y_{di}\mid f^{*}_{d}\sim\mathcal{N}(f^{*}_{d}(x_{i}),\sigma_{d}^{2})$ (15)
An alternative way of specifying $C$ is through a set of input-dependent covariance functions (Gneiting et al., , 2010; Bourotte et al., , 2016). In our study, we however only consider an input-independent $C$, as modeling the across-dimension correlations is not the main focus of this paper. In Section 6, we provide a more detailed discussion on this topic.
@@ 3.3.2 Latent variable HSGPs
Within latent-variable GPs, the inputs $x$ are considered as unobserved and are treated similar to other estimable parameters. To that end, we specify an observed quantity $\tilde{x}$ that guides the estimation of the latent inputs $x$. From a Bayesian perspective, $\tilde{x}$ acts as a prior-like object for the latent $x$ to be then further refined by the GPs learning from $y$. Specifically, we consider the scenario where $\tilde{x}$ is a noisy measurement of $x$. If we assume that the measurements $\tilde{x}$ are Gaussian with known measurement SD $s$, we can write for each observation $i$:
@@ 3.4 Bayesian inference
$p(y,x,\theta\mid\tilde{x})=p(x\mid\tilde{x})\prod_{d}^{D}p(y_{d}\mid x,\theta_{d})\,\,p(\theta_{d}).$ (18)
@@
$p(x,\theta\mid y,\tilde{x})=\frac{p(y,x,\theta\mid\tilde{x})}{\int\int p(y,x,\theta\mid\tilde{x})\,dx\,d\theta}.$ (19)
@@ 4 Simulation Study
For real-world data, we lack the ground truth values of latent variables. In contrast, in a simulated setting, we have full control over the data generating process including the true latent values. Thus, it is crucial to investigate the statistical properties of latent variable models using simulated data. To this end, we investigate exact GPs, HSGPs and VIGPs under various simulation scenarios exploring their behavior in terms of model convergence, model calibration, and estimation accuracy. The data generating conditions are inspired by Mukherjee et al., (2025).
@@ 4.1 Data generating scenarios
In our simulations, we consider two version of this process, one with low oscillations (high true $\rho_{d}$ values) and one with higher oscillations (low true $\rho_{d}$ values). These two periodic data scenarios mimic the challenges of modeling starkly different functional smoothness with the higher oscillations being the more challenging one. The hyperparameters of the periodic process closely resemble that of the GP data scenarios and are thus named the same for simplicity. The sampling distributions are also similarly specified as mentioned previously. For the periodic data with higher oscillation, we specify $\rho\sim\text{Normal}^{+}(0.5,0.05^{2})$ as compared to the lower oscillation case where $\rho\sim\text{Normal}^{+}(1,0.25^{2})$.
@@ 4.2 Model specifications
We fit exact GPs, HSGPs, and VIGPs for all simulation scenarios, with the exception of omitting exact GPs for sample sizes higher than $N=20$ due to the scalability issues discussed earlier (see Section 1). For the HSGPs, we select the boundary conditions $L$ by multiplying a scaler adjustment $c=1.25$ to the range of the input prior $\tilde{x}$. This way, we ensure that we have a compact cover [L,L][-L,L] that prevents input values $x$ to be near the boundaries. By doing so, we avoid model convergence issues completely (see Section 4.3). We select the minimum number of basis functions $M_{\rm min}$ loosely based on the suggestions of Riutort-Mayol et al., (2022). The choices for $M_{\rm min}$ depends on the choice of covariance function to accurately approximate the degree of non-linearity. Thus, for the SE covariance function,
@@
Both exact GPs and HSGPs are implemented in Stan (Stan Development Team, , 2024) and fitted with a single MCMC chain of 2000 iterations of which the first 1000 are discarded as warm-up. As shown in Mukherjee et al., (2025), model convergence is similar for multiple chains when fitting latent variable GPs. Thus, we run only a single chain per model to reduce overall computation times. The simulation studies are conducted on 50 vCPUs (Intel(R) Xeon(R) Gold 6230R CPU @ 2.10 GHz) with 720 GB work memory. The average runtime per dataset for the HSGPs with $N=200$ and $D=20$, the most computationally expensive case, was approximately 28 minutes for the SE model, 42 minutes for the Matern 3/2 model, and 34 minutes for the Matern 5/2 model. In the same scenarios, the VIGPs took approximately 2 minutes regardless of the covariance function. The exact GPs (just for $N=20$ and $D=20$) took 48 minutes for the SE model, 1.6 hours for the Matern 3/2 and 2.4 hours for the Matern 5/2 model. The HSGPs were significantly faster than the exact GPs but slower than VIGPs. The slower speed of HSGPs compared to VIGPs is counter-balanced by highly consistent model calibrations, superior estimation accuracy for latent variables, and overall much more stable model performance, as shown below.
@@ 4.3 Model convergence
In Figure 1, we show $\widehat{R}$, Bulk-ESS, and Tail-ESS for the latent $x$ and GP hyperparameters of both exact GPs and HSGPs with SE covariance function. While the exact GPs reach and exceed the relaxed $\widehat{R}$ threshold of 1.1 for some simulated datasets, the HSGPs consistently satisfy the much stricter 1.01 threshold of model convergence. HSGPs subsequently also have much higher Bulk and Tail-ESS. Overall, based on the diagnostics, HSGPs show much more consistent and stable convergence as compared to exact GPs. We provide convergence diagnostic figures for the other simulation scenarios in the supplementary materials Section C (Figures S3-S7)where we see similar results.
@@ 4.4 Testing model calibration
To that end, briefly, SBC aims to test the goodness of posterior approximations by exploiting self-consistency properties of Bayesian models. This results in a rank statistic that is uniformly distributed under the assumption of a well calibrated model (Modrák et al., , 2023), which can subsequently be checked both graphically and numerically. As a graphical test, we plot the empirical cumulative distribution function (ECDF) of the ranks along with their 95% confidence regions under the assumption of uniformity (Säilynoja et al., , 2022). ECDFs lying outside of their confidence region indicate miscalibrations. These graphical tests also yield a test statistic (the log $\gamma$ score; see Säilynoja et al., , 2022), which we further analyze numerically. More details on SBC is provided in supplementary materials Section A.
Figure 3: Squared exponential scenario: log $\gamma$ scores offset by the 95% confidence threshold for all the fitted models. The behavior of scores across true latent $x$ values are shown in the right-hand panel. The blue dashed line denotes the threshold to reject uniformity. The HSGP($M$) shows the HSGPs with their corresponding number of basis functions.
@@
To streamline calibration checking across simulation scenarios, we analyze the log $\gamma$ scores using a multi-level model (Bürkner, , 2017) described in supplementary materials Section A. In Figure 3, we show the predicted log $\gamma$ scores of the different GP models under the SE simulation scenario.
Figure 4: Squared exponential scenario (highly varying $\rho$): log $\gamma$ scores offset by the 95% confidence threshold for all the fitted models. The behavior of scores across true latent $x$ values are shown in the right-hand panel. The blue dashed line denotes the threshold to reject uniformity. The HSGP($M$) shows the HSGPs with their corresponding number of basis functions.
For all the choices of $N$ and $D$, the latent variable estimates from HSGPs are consistently well calibrated. By comparison, the exact GPs fail the calibration test for $D=20$. The VIGPs shows miscalibrations especially for higher output dimension cases throughout all the choices of $N$. Additionally, exact GPs exhibit worse calibrations (decreasing log $\gamma$ scores) when estimating larger values of the latent inputs. This issue directly causes problems in estimating latent values closer to the boundary. VIGPs show similar behavior for lower sample sizes $N=20$ although not as strong as exact GPs. The HSGPs are able to overcome and rectify this behavior. For the special scenario of SE data with highly variable length-scales as shown in Figure 4, VIGPs and HSGPs with small number of basis function struggle to pass the calibration tests, especially for $N=200$. The miscalibrations are however alleviated with a higher number of basis functions in HSGPs. The log $\gamma$ results for the Matern 3/2 and 5/2 simulation scenarios are presented in supplementary materials Section D (Figures S8 and S9).